A 6-quart cooling system is checked and found to be filled with a solution that is 20% antifreeze. The desired strength of the solution is 60 % antifreeze. How many quarts of solution need to be drained and replaced with pure antifreeze to reach the desired strength?
Answer by funmath(2933)

A 6-quart cooling system is checked and found to be filled with a solution that is 20% antifreeze. The desired strength of the solution is 60 % antifreeze. How many quarts of solution need to be drained and replaced with pure antifreeze to reach the desired strength?
You start with 6 quarts of 20% antifreeze .20*(6)=1.2
You take out (x amount) -.20x=-.2x
You put in 100% of (the same x amount) +1.00x=+x
You want 6 quarts of 60% antifreeze .60(6)=3.6
The problem to solve:
.20(6)-.20x+1.00x=.60(6)
1.2-.2x+x=3.6
1.2+.8x=3.6
1.2-1.2+.8x=3.6-1.2
.8x=2.4
.8x/.8=2.4/.8
x=3 qts
Happy Calculating!!!